OpenStudy (jellybot23):

Literally no idea and have been on it for an hour! Prove: (tanx+secx)^2 = 2 secx+2tanxsecx-1

OpenStudy (anonymous):

square it !!

OpenStudy (jellybot23):

Square the right hand side?

OpenStudy (anonymous):

then use a well known identity involving secant squared and tangent squared

OpenStudy (anonymous):

no, the left hand side has an exponent of two multiply that out

OpenStudy (jellybot23):

Well I get tan^2x+2tanxsecx+sec^2x ?

OpenStudy (jellybot23):

@satellite73

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now look for an identity involving secant squared and tangent square you are stuck with the \(2\tan(x)\sec(x)\)

OpenStudy (jellybot23):

I know that sec^2x-1=tan^2x

OpenStudy (jellybot23):

I just am not the best with trigonometry, so I get confused with this sometimes

OpenStudy (anonymous):

what happesn if you replace \(\tan^2(x)\) by \(\sec^2(x)-1\)?

OpenStudy (jellybot23):

Would the secants cancel out and I would be left with 2tanxsecx -1?

OpenStudy (jellybot23):

Oh wait, no, the secants would add together!

OpenStudy (jellybot23):

Ohhhhh!

OpenStudy (anonymous):

yeah, but you don't actually get what you wrote above, you get \[2\sec^2(x)+2\sec(x)\tan(x)-1\]

OpenStudy (jellybot23):

That's the right side of the equation, though! So that's all we need to do?

OpenStudy (anonymous):

it is ? you didn't put a square over the secant, but if it i supposed to be there then we are done

OpenStudy (anonymous):

in fact it has to be there, so maybe it was just a typo

OpenStudy (jellybot23):

Oh sorry, yes... I am looking at my paper and it is (tanx+secx)^2 = 2 sec^2x+2tanxsecx-1

OpenStudy (jellybot23):

Awesome! Thanks so much :)

OpenStudy (anonymous):

yw now too bad was it (certainly not worth an hour of your time!)

OpenStudy (jellybot23):

Haha not at all! Some proofs come really easy to me and others for some reason stump me! Lol; I think I overthink it :p